A friend throws a baseball horizontally. He releases it at a height of 2.0 m and it lands 21 m from his front foot, which is directly below the point at which he released the baseball. (a) How long wa
1
Expert's answer
2012-09-18T11:05:28-0400
a) t=sqrt(2(2+21)/g)=sqrt(4.6)=2.2